OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? 7(4x + 10) + 3(8 - x) A.31x + 94 B.25x + 46 C.25x - 94 D.25x + 94

OpenStudy (cggurumanjunath):

7(4x+10) + 3(8-x) ; 7(4x+10) =28x+70 ; 3(8-x) = 24 -3x ; 28x+70 +24 -3x =25x +94

OpenStudy (cggurumanjunath):

use distributive property

OpenStudy (cggurumanjunath):

a(b+c) =ab+ac